A gardener pushes a lawnmower with a force of 1000 N directed along its handle which makes an angle of 60º with the ground in order for it to move at a constant velocity. You are required to calculate the following:

a. The component of this force which is directed horizontally, H, in order to overcome the frictional forces and maintain a constant velocity.

b. the component of the force which is directed vertically downwards, V .


1
Expert's answer
2021-02-28T07:22:21-0500
"F_H=F\cdot cos\theta=1000\ N\cdot cos60^{\circ}=500\ N,""F_V=F\cdot sin\theta=1000\ N\cdot sin60^{\circ}=866\ N."